Required information Exercise 10-3A Record the issuance of common stock (LO10-2) [The following information applies to the quView transaction list Journal entry worksheet < 1 2 Record the issuance of 600 shares of common stock for $40 per share. NoteJournal entry worksheet < 1 2 Record the issuance of 100 additional shares of common stock for $44 per share. Note: Enter deb2. Record the transactions, assuming Clothing Frontiers has either $1 par value or $1 stated value common stock. (If no entry2. Record the transactions, assuming Clothing Frontiers has either $1 par value or $1 stated value common stock. (If no entry
Answer #1
1
Debit Credit
January 1 Cash 24000 =600*40
     Common stock 24000
April 1 Cash 4400 =100*44
     Common stock 4400
2
Debit Credit
January 1 Cash 24000 =600*40
     Common stock 600
     Additional paid­-in capital 23400
April 1 Cash 4400 =100*44
     Common stock 100
     Additional paid­-in capital 4300